

********************************************************************
UARS WEEKLY STATUS REPORT - 23 March 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
SUSIM performed its daily solar scans this past week during
periods of predicted UARS telemetry.
We have made a request to the Project and fellow SSPP instrument teams
that the SSPP be operated in closed loop tracking mode during our
upcoming major calibration slated for early May. The improved tracking
would improve the reliability of the calibration.
The calibrations necessary for version 21 of the SUSIM data are complete.
We expect to begin reprocessing when the CDHF rehosting is finished.
***********
ACRIM
***********
The Instrument Science Team continues re-processing ACRIM II data. Target
date for release of version 2 results is April or May, 2001. The ACRIM II
instrument is operating normally, obtaining a full set of data each day.
No significant impact of the SSPP pointing degradation has been detected.
***********
PEM
***********
PEM/AXIS was activated in its half power mode on March 19 and deactivated
on March 13. At the current time, AXIS is not generating energy deposition.
SwRI is looking into this. Tests of AXIS production code on the rehost computer
seem to indicate that there are system file access violations on the rehost
computer, which are as yet unexplained.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
PEM has examined its processed files of the rehost machine and has found
differences between those and the cataloged files in the UARS database. The
PEM investigation found that the Level 0 files used to generate the rehost
data were not the same as those used in previous production. Therefore, PEM
finds that the test is not valid and does NOT APPROVE production of PEM data
on the rehost machine at this time.
The PEM team is working on replacing the PEM:: machine with a DEC Alpha
machine to be called PEM2K::. PEM has requested DAP command capability with
the CMS from PEM2K::. Tests of the NRT data transfer from CDHF2:: through
PEM2K:: to the unix computer running the PEM GSE software show no difference
between equivalent NRT files transferred from PEM::. Thus, NRT data from
CDHF2:: is captured correctly on PEM2K::. Tests of the NRT data transfer
from CDHFB:: were not successful from both PEM:: and PEM2K::. Results of the
test of NRT data transfers from the rehost computer acted like the rehost
computer did not recognize either PEM machine. Investigation is underway.
PEM has been working with GSFC on allowing commanding of the PEM
instrument by creating DAPs on the PEM2K machine and submitting them to the
CMS. So far, this has not been achieved. It was discovered that PEM2K is
running Openwindows version 7.2-1 which operates differently across the
network than Openwindows version 7.1-2 (the CDHFB operating system). Compaq
is investigating to determine a solution.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI continues to function normally, making wind measurements during
the sunlit parts of orbits nearly every day, and measuring nighttime winds
when power is available.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Daily processing of instrument Daily Activity Plans continues
with no problems. Instrument operations schedule is now complete
to 19 April 2001.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 10 March 2001 (Orbit 51901, GMT Day 069) through 16 March
2001 (Orbit 52005, GMT Day 075).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 070/2256 51930 Power OFF
PEM 071/2038 51943 Calibration
PEM 072/0041 51946 PEM to day only
PEM 072/2300 51960 AXIS 1 OFF
PEM 073/2302 51975 Power OFF
ACRIM 073/2302 51975 Power OFF
SUSIM 073/2302 51975 Power OFF
SOLSTICE 074/0004 51975 Day only operations
HRDI 074/2235 51989 INST PB OFF
A Forward-to-Reverse Yaw Around maneuver (number 97) was successfully executed
in Orbit 51984 beginning 074/15:16:52 and completing at 074/15:48:00 on 15
March 2001. The SOLSTICE instrument was in day only and all other instruments
were in their off or standby mode (including HRDI and PEM). Due to the SA
position at 269 degrees, the effective night length of the spacecraft after the
maneuver was approximately 55.4 minutes. Post maneuver analysis of data
indicate that during the Yaw Around, the end-of-night load bus voltage was
24.10 volts and the average depths of discharge (DODs) on Batteries 2 and 3
were approximately 22.0 and 23.0 percent respectively. The next (reverse-to-
forward) YAW is scheduled for 19 April 2001 (DOY 109).
The beta angle ranged between the angles of 21.4 degrees to 01.5 degrees for
this report period. The beta angle is now decreasing towards a minimum angle
of 52 degrees on 02 April 2001 (DOY 074).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.2 and 1.3 degrees C. Additional battery performance data for each
battery on 09 March 2001 and 16 March 2001 are provided in Appendix D.
The clock error ranged from -10.4 to 11.9 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 11
¥ PEM = 09
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 01
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 02/026 - 2287. 49920005
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/067 show the observatory data loss to be 300 hours, 20
minutes, 29.464 seconds. This shows an overall increase of 29.616 seconds since
the last report period. This is a 0.39 percent data loss, which equals a 99.61
percent data capture for the mission.
The increase of 29.616 seconds occurred over two (2) days:
· 15.072 seconds on Day 071 due to No Data Overlap
· 14.544 seconds on Day 074 due to No Data Overlap
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UARS is using both the Fine Sun Sensor (FSS) and the Three Axis Magnetometer
(TAM) as inputs to UFLTR this week. The FOT is pleased with their performance,
but continues to make adjustments to the TAM alignment to improve its
performance and UARS' attitude. The FOT awaits the completion of the FSW patch
that will improve the magnetic field model from a 5th order to 7th order
solution.
MEETINGS
The FOT attended a TAC Replacement Meeting on Tuesday, March 13, 2001.
The FOT attended a TAC Replacement Meeting on Friday, March 16, 2001
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 19 April 2001 (DOY 109). This maneuver
will be a Reverse to Forward Flight maneuver (Southern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 16 March 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
51912 069/1845 01-150 NBTR NOOPS D070
51920 070/0726 01-150 NBTR NOOPS D-070
51930 070/2256 01-151 HALOE Power OFF
51941 071/1741 01-152 OBC MAGMOD Time Update
51943 071/2038 01-153 PEM Calibration
51944 071/2114 01-154 ACS TAM/FSS Monitor/Update Mode
51944 071/2114 01-155 ACS TAM Alignment
51946 072/0041 01-153 PEM PEM to day only
51960 072/2300 01-156 PEM AXIS 1 OFF
51969 073/1423 01-157 OBC Clk Adj fr 0.0 to -20.6 msec/day
51975 073/2302 01-158 PEM Power OFF
51975 073/2302 01-159 ACRIM Power OFF
51975 073/2302 01-160 SUSIM Power OFF
51975 074/0004 01-161 SOLSTICE Day only operations
51984 074/1424 01-162 ACS TAM to MONITOR MODE
51984 074/1424 01-163 ACS YAW MANEUVER FWD to REV
51985 074/1619 01-164 ACS TAM ALIGN/ UPDATE MODE
51987 074/1911 01-166 ACS TAM GAIN CHANGE
51989 074/2235 01-167 HRDI INST PB OFF
51986 074/1723 01-165 OBC Clk Adj fr -20.6 to 0.0 msec/day
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 16 March 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period 51883
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 16 March 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
51969 073/1429 Clk Adj fr 0.0 to -20.6 msec/day
51986 074/1734 Clk Adj fr -20.6 to 0.0 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 16 March 2001
09 Mar 2001 - GMT Day 068 Beta = 25.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- --------- ------ ------- -----
1 +016.8/+00.0 -3.05/-4. +0.0/+0.0 13.0 0.000 0.0
2 +16.8/-89.6 +5.79/+4.51 +30.0/-14.4 24.5 1.030 21.1
3 +22.4/-95.2 +4.51/+3.55 +29.6/-14.4 24.3 1.020 20.9
16 Mar 2001 - GMT Day 075 Beta = 01.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- --------- ------ ------- -----
1 +022.4/+00.0 -3.68/-4.94 +0.0/+0.0 13.0 0.000 0.0
2 +16.8/-89.6 +5.15/+4.19 +29.6/-13.6 25.3 1.020 20.4
3 +33.6/-44.8 +3.87/+2.92 +28.8/-13.2 25.3 1.020 20.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
---- ---- ---------
-no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 16 March 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-014 FOT 51921 070/0909 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-014 FOT see list Mar 01 14170 Generic Late Acquisition
SEVNTFW01/11